OpenStudy (anonymous):

Solve the following equations, giving your ans correct to 3 sig. fig. thx u. a) 2e^(2x+1) = e^(x+1) +15e b) 2e^2x = 5e^(x+1) - 2e^2

OpenStudy (amistre64):

so how would you go about solving this?

OpenStudy (amistre64):

this property of exponents might come in handy:\[\large b^{n+m}=b^n*b^m\]

OpenStudy (zarkon):

u=e^x for the 2nd

OpenStudy (anonymous):

first one, divide both sides by "e" to get \[2e^{2x}=e^x+15\] then perhaps \[2e^{2x}-e^x-15=0\] a nice quadratic equation, that you can solve by factoring

OpenStudy (anonymous):

second one \[2e^{2x} = 5e^{x+1} - 2e^2\] \[2e^{2x}=5e(e^x)-2e^2\] \[2e^{2x}-5e(e^x)+2e^2=0\] quadratic formula for this one